# Cartesian Forest Matching

### v1

@@ Abstract
In this paper, we introduce the notion of Cartesian Forest, which generalizes Cartesian Trees, in order to deal with partially ordered sequences. We show that algorithms that solve both exact and approximate Cartesian Tree Matching can be adapted to solve Cartesian Forest Matching in average linear time. We adapt the notion of Cartesian Tree Signature to Cartesian Forests and show how filters can be used to experimentally improve the algorithm for the exact matching. We also show a one to one correspondence between Cartesian Forests and Schröder Trees.
Keywords:
Cartesian Tree Cartesian Forest Pattern Matching Approximate Pattern Matching Schröder Tree
@@ 1 Introduction
Indexing structures in the Cartesian Tree pattern matching framework are presented in [12, 10, 14]. Methods for computing regularities are given in [9] and methods for computing palindromic structures are presented in [6]. An algorithm for episode matching (given two sequences $p$ and $t$, finding all minimal length factors of $t$ that contains $p$ as a subsequence) in Cartesian Tree framework is presented in [13]. Practical methods for finding longest common Cartesian substrings of two strings appeared in [5]. Very recently, dynamic programming approaches for approximate Cartesian Tree pattern matching with edit distance has been considered in [11] and longest common Cartesian Tree subsequences are computed in [18]. Efficient algorithms for determining if two equal-length indeterminate strings match in the Cartesian Tree framework are given in [8].
Cartesian Trees are defined on arbitrary sequences where there can be values with multiple occurrences. In that case, ties can be broken in any way. In this article, in order to better deal with equal values we introduce Cartesian Forests.
We show that algorithms that solve both exact and approximate Cartesian Tree Matching can be adapted to solve Cartesian Forest Matching in average linear time. We then show a one to one correspondence between Cartesian Forests and Schröder Trees. More specifically we give algorithms for computing a Schröder Tree and a special type of words with parentheses given a Cartesian Forest. We also adapt the notion of Cartesian Tree Signature introduced in [4] to Cartesian Forests, and show how this notion can be used to experimentally improve the computation.
2 Definitions
In this paper, a sequence is always defined on an ordered alphabet. For a given sequence $x$, $|x|$ denotes the length of $x$. A sequence $v$ is factor of a sequence $x$ if $x=uvw$ for any sequences $u$ and $w$. A sequence $u$ is a prefix (resp. suffix) of a sequence $x$ if $x=uv$ (resp. $x=vu$). For a sequence $x$ of length $m$, $x[i]$ is the $i$-th element of $x$ and $x[i\ldots j]$ represents the factor of $x$ starting at the $i$-th element and ending at the $j$-th element, for $1\leq i\leq j\leq m$.
2.1 Cartesian Tree and Cartesian Forest
@@
Given a sequence $x$ of length $m$, the Cartesian Tree of $x$, denoted by $C(x)$, is the binary tree recursively defined as follows:
@@
  • if $x[1\ldots m]$ is not empty and $x[h]$ is the smallest value of $x$, $C(x)$ is the binary tree with $h$ as its root, the Cartesian Tree of $x[1\ldots h-1]$ as the left subtree and the Cartesian Tree of $x[h+1\ldots m]$ as the right subtree.
@@
We denote by $\operatorname{rb}(T)$ the list of nodes on the right branch of a Cartesian Tree $T$. The Cartesian Tree of a sequence can be built online in linear time and space [7]. Informally, let $x$ be a sequence such that $C(x[1\ldots h-1])$ is already known. In order to build $C(x[1\ldots h])$, one only needs to find the nodes $j_{1}<\cdots<j_{k}$ in $\operatorname{rb}(C(x[1\ldots h-1]))$ such that $x[j_{1}]>x[h]$. If $j_{1}$ is the root of $C(x[1\ldots h-1])$ then $h$ becomes the root of $C(x[1\ldots h])$ otherwise let $j_{0}$ be the parent node of $j_{1}$, then node $h$ will be the root of the right subtree of $j_{0}$. In both cases, $j_{1}$ will be the root of the left subtree of node $h$. Then $\operatorname{rb}(C(x[1\ldots h])=\operatorname{rb}(C(x[1\ldots h-1]))\setminus(j_{1},\ldots,j_{k})\cup(h)$. All these operations can be easily done by implementing $\operatorname{rb}$ with a stack. The amortized cost of such an operation can be shown to be constant.
@@
Cartesian Trees are defined on arbitrary sequences where there can be multiple occurrences of the smallest value.In that case, ties can be broken in any way. Usually the first occurrence of the smallest value is chosen to be the root of the tree. In order to better deal with equal values we now introduce Cartesian Forests. We first introduce the combinatorial object in itself, then we define what the Cartesian Forest of a sequence is.
Definition 2 (Cartesian Forest $F$)
A Cartesian Forest $F$ can be:
  • empty,
  • a sequence of $k$ planar trees rooted in $(r_{1},\ldots,r_{k})$, with $k\geq 1$, such that $r_{1}$ has a left Cartesian sub-Forest and a right Cartesian sub-Forest, and $r_{i}$ has only a right Cartesian sub-Forest for all $i\in\{2,\ldots,k\}$ (the left sub-Forest is necessarily empty).
Figure 1: On the left is a valid Cartesian Forest. In the middle, it is not a valid Cartesian Forest because the second tree in the right sub-Forest of $r^{\prime}_{1}$ has a left sub-Forest. On the right, a planar forest is not Cartesian: the second tree has a left sub-Forest.
Definition 3 (Cartesian Forest $F(x)$ of a sequence $x$)
  • if $x$ is empty, then $F(x)$ is the empty forest;
  • if $x[1\ldots m]$ is not empty, let $(r_{1},\ldots,r_{k})$ be the ordered sequence of all the $k$ positions of the smallest value of $x$, then $F(x)$ is a forest composed of $k$ planar trees whose roots are $r_{1},\ldots,r_{k}$, such that:
    • the Cartesian Forest $F(x[1\ldots r_{1}-1])$ forms a sequence of left subtrees of $r_{1}$,
    • for $1\leq i<k$, the Cartesian Forest $F(x[r_{i}+1\ldots r_{i+1}-1])$ forms a sequence of right subtrees of $r_{i}$,
    • the Cartesian Forest $F(x[r_{k}+1\ldots m])$ forms a sequence of right subtrees of $r_{k}$.
Figure 2 shows an example of the Cartesian Forest of a sequence.
Figure 2: Cartesian Forest associated to an ordered sequence $x$.
3 Exact Cartesian Forest Matching using Linear Representations
We adapt the linear representations of Cartesian Trees, such as the Parent-Distance [15] and the Skipped-Number representation [4] to Cartesian Forests. Basically, for a sequence $x$, when computing $C(x[1\ldots h])$ from $C(x[1\ldots h-1])$, the Parent-Distance of $h$ is equal to the distance between $h$ and its parent in $C(x[1\ldots h])$ and its Skipped-Number is the number of nodes removed from the right branch of $C(x[1\ldots h-1])$ comparing to the right branch of $C(x[1\ldots h])$ (see [1]).
The main idea of the adaptation to Cartesian Forests is simple: both representations involve the comparisons of two values $x[i]$ and $x[j]$ in the original sequence $x$, for a given $i<j$. Suppose we have either $x[i]>x[j]$ or $x[i]<x[j]$, then we are in the case of Cartesian Tree and the linear representation is the same.
The idea for the Parent-Distance of the Cartesian Forest of a sequence $x$ is that its absolute value at a position $h$ gives the distance between a node $h$ and its parent or left sibling in the Cartesian Forest associated to the sequence $x[1\ldots h]$.
Formally, let $\operatorname{small}_{x}(h)=\max_{1\leq j<h}\{j\ |\ x[j]<x[h]\}\cup\{0\}$ and $\operatorname{equal}_{x}(h)=\max_{1\leq j<h}\{j\ |\ x[j]=x[h]\}\cup\{0\}$.
Definition 4 (Parent-Distance representation $PD_{x}$)
@@ 3 Exact Cartesian Forest Matching using Linear Representations
Given a sequence $x[1\ldots m]$, the Parent-Distance representation of $x$ is an integer sequence $PD_{x}[1\ldots m]$ defined as follows:
$PD_{x}[h]=\begin{cases}h-\operatorname{small}_{x}(h)&\text{if }\operatorname{small}_{x}(h)>\operatorname{equal}_{x}(h)\cr-(h-\operatorname{equal}_{x}(h))&\text{if }\operatorname{small}_{x}(h)<\operatorname{equal}_{x}(h)\cr 0&\text{otherwise.}\end{cases}$
Definition 5 (The referent table $\operatorname{ref}_{x}$)
Given a sequence $x[1\ldots m]$, the referent table of $x$ is a sequence of sets $\operatorname{ref}_{x}[1\ldots m]$ such that
$\operatorname{ref}_{x}[h]=\begin{cases}\min_{h<j\leq m}\{j\ |\ x[j]\leq x[h]\}&\text{if such }j\text{ exists}\cr-1&\text{ otherwise.}\end{cases}$
@@
Given a sequence $x[1\ldots m]$, the Skipped-number representationof $x$ is an integer sequence $SN_{x}[1\ldots m]$ such that
$SN_{x}[h]=\begin{cases}|\{j<h\mid\operatorname{ref}_{x}[j]=h\}|&\mbox{if }\operatorname{small}_{x}(h)>\operatorname{equal}_{x}(h)\cr-|\{j<h\mid\operatorname{ref}_{x}[j]=h\}|&\mbox{if }\operatorname{small}_{x}(h)<\operatorname{equal}_{x}(h)\cr\end{cases}$
Figure 3 shows examples of Parent-Distance representations, Skipped-number representations and referent tables of two Cartesian Forests.
Figure 3: Two sequences $x$ and $y$, and their associated Cartesian Forests $F(x)$ and $F(y)$ and their corresponding Parent-Distance representations, Skipped-number representation and referent tables. As one can see, $x$ is a prefix of $y$ and the forest $F(x)$ is transformed into a sequence of left subtrees in $F(y)$.
Definition 7 (Cartesian Forest Matching (CFM))
Given two sequences $p[1\ldots m]$ and $t[1\ldots n]$, find every position $j$, with $11\leq j\leq n-m+$ 1, such that $t[j\ldots j+m-1]\approx_{CF}p[1\ldots m]$.
Example 1
Let $t=(5,7,3,6,3,7,2,8,2,4,3,3)$ and $p=(2,3,1,4,1,5)$ respectively be the text and the pattern. We have two occurrences of $p$ in $t$ as
$t[1\ldots 6]=(5,7,3,6,3,7)\approx_{CF}p\text{ and }t[5\ldots 10]=(3,7,2,8,2,4)\approx_{CF}p.$
@@ 2 $x\leftarrow t[1\ldots m]$
3 $LN_{p},LN_{x}\leftarrow$ linear representations of $C(p)$ and $C(x)$

### v2

@@ Abstract
In this paper, we introduce the notion of Cartesian Forest, which generalizes Cartesian Trees, in order to deal with partially ordered sequences. We show that algorithms that solve both exact and approximate Cartesian Tree Matching can be adapted to solve Cartesian Forest Matching in linear time in the worst-case for exact matching and linear time on average for approximate matching. We adapt the notion of Cartesian Tree Signature to Cartesian Forests. We also show a one-to-one correspondence between Cartesian Forests and Schröder Trees.
@@ 1 Introduction
Indexing structures in the Cartesian Tree pattern matching framework are presented in [16, 13, 18]. Methods for computing regularities are given in [12] and methods for computing palindromic structures are presented in [9]. An algorithm for episode matching (given two sequences $p$ and $t$, finding all minimal length factors of $t$ that contains $p$ as a subsequence) in Cartesian Tree framework is presented in [17]. Practical methods for finding longest common Cartesian substrings of two strings appeared in [7]. Very recently, dynamic programming approaches for approximate Cartesian Tree pattern matching with edit distance have been considered in [14] and longest common Cartesian Tree subsequences are computed in [22]. Efficient algorithms for determining if two equal-length indeterminate strings match in the Cartesian Tree framework are given in [11].
Cartesian Trees are defined (Section 2) on arbitrary sequences where there can be values with multiple occurrences. In that case, ties should be broken in some way. In this article, in order to better deal with equal values we introduce Cartesian Forests (Section 3).
We show that algorithms that solve both exact and approximate Cartesian Tree Matching (Section 4 and Section 5) can be adapted to solve Cartesian Forest Matching in linear time in the worst-case for exact matching and linear time on average for approximate matching. We then show a one to one correspondence between Cartesian Forests and Schröder Trees (Section 6). More specifically we give algorithms for computing a Schröder Tree and a special type of words with parentheses given a Cartesian Forest. We also adapt the notion of Cartesian Tree Signature introduced in [6] to Cartesian Forests (Section 7), and show how this notion can be used to experimentally improve the computation (Section 8).
2 Preliminaries
In this paper, a sequence is always defined on an ordered alphabet. For a given sequence $x$, $|x|$ denotes the length of $x$. A sequence $v$ is factor of a sequence $x$ if $x=uvw$ for any sequences $u$ and $w$. A sequence $u$ is a prefix (resp. suffix) of a sequence $x$ if $x=uv$ (resp. $x=vu$). A prefix of a sequence that is also a suffix is called a border. For a sequence $x$ of length $m$, $x[i]$ is the $i$-th element of $x$ and $x[i\ldots j]$ represents the factor of $x$ starting at the $i$-th element and ending at the $j$-th element, for $1\leq i\leq j\leq m$.
2.1 Cartesian Trees
@@
Given a totally ordered sequence $x$ of length $m$, the Cartesian Tree of $x$, denoted by $C(x)$, is the binary tree recursively defined as follows:
@@
  • if $x[1\ldots m]$ is not empty and $x[i]$ is the smallest value of $x$, $C(x)$ is the binary tree with $i$ as its root, the Cartesian Tree of $x[1\ldots i-1]$ as the left subtree and the Cartesian Tree of $x[i+1\ldots m]$ as the right subtree.
@@
We denote by $\operatorname{rb}(T)$ the list of nodes on the right branch of a Cartesian Tree $T$. Also, $C_{i}(x)=C(x[1\ldots i])$ to simplify the notations. The Cartesian Tree of a sequence can be built online in linear time and space [10]. Informally, let $x$ be a sequence such that $C_{i-1}(x)$ is already known. In order to build $C_{i}(x)$, one only needs to find the nodes $j_{1}<\cdots<j_{k}$ in $\operatorname{rb}(C_{i-1}(x))$ such that $x[j_{1}]>x[i]$. Then, $j_{1}$ will be the root of the left subtree of node $i$. If $j_{1}$ is the root of $C_{i-1}(x)$ then $i$ becomes the root of $C_{i}(x)$, otherwise let $j_{0}$ be the parent node of $j_{1}$, then node $i$ will be the root of the right subtree of $j_{0}$. Then $\operatorname{rb}(C_{i}(x))=\operatorname{rb}(C_{i-1}(x))\setminus(j_{1},\ldots,j_{k})\cup(i)$. All these operations can be easily done by implementing $\operatorname{rb}$ with a stack. The amortized cost of such an operation can be shown to be constant.
@@
Cartesian Trees are defined on arbitrary sequences where there can be multiple occurrences of the smallest value. In that case, ties should be broken in some way. Usually the first occurrence of the smallest value is chosen to be the root of the tree.
Given a Cartesian Tree $C(x)$ of a sequence $x$ of length $m$, a linear representation of $C(x)$, denoted $\operatorname{LN}_{x}$, is a table of integers of length $m$, and the set of linear representations of a given length is in bijection with the set of Cartesian Trees of the same size. A linear representation of Cartesian Trees was introduced by Park et al. [19]: the parent-distance representation (see example Figure 3). Using this representation and classical string algorithms (Knuth-Morris-Pratt (KMP) [15] and Aho-Corasick [1] algorithms), they obtained linear-time solutions for single and multiple pattern Cartesian Tree matching (see Figure 1 for an example).
Let $\operatorname{small}_{x}(i)=\max(\{k\mid x[k]<x[i]\mbox{ with }1\leq k<i\}\cup\{0\})$.
Definition 2(Parent-distance representation $\operatorname{PD}_{x}$)
@@
Given a sequence $x$ of length $m$, the parent-distance representation of $x$ is an integer sequence $\operatorname{PD}_{x}[1\ldots m]$, which is defined as follows:
$\operatorname{PD}_{x}[i]=\begin{cases}i-\operatorname{small}_{x}(i)&\text{if }\operatorname{small}_{x}(i)>0\cr 0&\text{otherwise.}\end{cases}$
Since the parent-distance representation has a one-to-one mapping with Cartesian Trees, it can replace them without loss of information.
The parent-distance table is not the only linear representation of Cartesian Trees. Another linear representation is based on the number of nodes skipped by a new node when building online the Cartesian Tree of a sequence. Informally, for each node $i$, it stores the number of nodes $(j_{1}<\cdots<j_{k})$ deleted from the right path when computing $C_{i}(x)$ from $C_{i-1}(x)$. We say that node $i$ skipped nodes $(j_{1}<\cdots<j_{k})$ and that these nodes are skipped by node $i$. The left subtree of a Cartesian Tree $C(x)$ is denoted $left(C(x))$ and the length of its right branch $\operatorname{RB}(C(x))=|\operatorname{rb}(C(x))|$.
See Figure 1 for an example.
@@
Given a sequence $x[1\ldots m]$, the Skipped-number representation of $x$ is a sequence $SN_{x}[1\ldots m]$ of integers such that $SN_{x}[i]$ is the length of the right-branch of the left-subtree of the subtree rooted at node $i$ of the Cartesian Tree of $x$, that is $SN_{x}[i]=\operatorname{RB}(\operatorname{left}(C_{i}(x)))$.
Both linear representations can be computed in linear time [19, 6, 3].
2.2 Exact Cartesian Tree Matching
Given two sequences $x$ and $y$ of length $m$ and $n$ respectively, the Exact Cartesian Tree Matching problem (CTM) consists of finding all the factors of $y$ that have the same Cartesian Tree of the one of $x$. This can be done in time $O(m+n)$ by using a KMP-like algorithm. To achieve this complexity, a Cartesian border table is used. Given the Parent-Distance representation of $x$, the Parent-Distance of a factor $x[i\ldots j]$ of $x$ satisfies:
$\operatorname{PD}_{x[i\ldots j]}[k]=\begin{cases}0&\mbox{if }\operatorname{PD}_{x}[i+k-1]\geq k,\cr\operatorname{PD}_{x}[i+k-1]&\mbox{otherwise}.\end{cases}$
Definition 4(Cartesian Border Table $CTBord$)
The Cartesian border table of $x$ is defined in the following way: $CTBord[0]=0$ and, for $1\leq k<i$, $CTBord[i]=\max\{k\mid C_{k}(x)=C(x[i-k+1\ldots i])\}$.
Figure 1 shows the Cartesian border table of $x=(3,1,6,4,8,6,7,5,9$).
$i$ 1 2 3 4 5 6 7 8 9
$x[i]$ 3 1 6 4 8 6 7 5 9
$\operatorname{PD}_{x}[i]$ 0 0 1 2 1 2 1 4 1
$SN_{x}[i]$ 0 1 0 1 0 1 0 2 0
$CTBord[i]$ 0 1 1 2 3 4 5 2 3
Figure 1: The parent-distance representation, the skipped-number representation and the Cartesian border table of a sequence $x$.
2.3 Approximate Cartesian Tree Matching
Given two sequences $x$ and $y$ of length $m$ and $n$ respectively, the Approximate Cartesian Tree Matching problem (ACTM) consists of finding all the factors of $y$ that have the same Cartesian Tree of the one of $x$ up to some differences.
Unfortunately, the linear-time solutions obtained in [19] for the exact case do not work anymore in the general approximate case. Also we conjecture that there is no algorithm that solves approximate pattern matching with a linear worst-case complexity.
In [3], the authors provide the following MetaAlgorithm that can solve ACTM with up to one difference (substitution, insertion, deletion, swap) in quadratic worst-case time complexity and in linear time in average. The test at line $5$, using the $\sim$ symbol, signifies that depending on the considered notion of error, we will not test equality between linear representation but rather test an equivalence.
@@ 2 $x\leftarrow t[1\ldots m]$
3 $\operatorname{LN}_{p},\operatorname{LN}_{x}\leftarrow$ linear representations of $C(p)$ and $C(x)$
@@ 8   Update $\operatorname{LN}_{x}$
3 Cartesian Forests
A limitation of Cartesian Trees is that it does not take into account equal values in the original sequence. For instance, the sequence $(1,1,1,1,1)$ shares the same Cartesian Tree as $(1,2,3,4,5)$ when ties are broken by ordering equal values according to their positions. In order to better deal with equal values we now introduce Cartesian Forests.
Recall that a planar tree, or general tree, is a tree in which each node possesses a sequence of children, or subtrees. In the following, trees designates planar trees.
Informally, during the construction of $C_{i}(x)$ from $C_{i-1}(x)$, assume that node $i$ should be inserted in the right branch between nodes $j_{0}$ and $j_{1}$ such that $x[j_{0}]\leq x[i]<x[j_{1}]$. In a Cartesian Tree, node $i$ becomes the root of the right subtree of node $j_{0}$ and node $j_{1}$ becomes the root of the left subtree of node $i$. In a Cartesian Forest, node $i$ becomes the root of the right subtree of node $j_{0}$ only if $x[j_{0}]<x[i]$ while it becomes the right sibling of node $j_{0}$ if $x[j_{0}]=x[i]$ and node $j_{1}$ remains in the right subtree of node $j_{0}$.
A forest is a sequence of trees. An empty forest is an empty sequence that does not contain any tree. The children of a node can be seen as a sub-forest. In the following, a node of a tree in a Cartesian Forest, instead of having a sequence of children, can possess a left sequence of children and a right sequence of children, which we call the left sub-forest and the right sub-forest. The leftmost tree of a Cartesian Forest is actually the only one that can possess a left sub-forest and a right sub-forest, whereas all the other trees can only have a right sub-forest. Figure 2 illustrates this with an example of a Cartesian Forest.
Definition 5(Cartesian Forest $F(x)$ of a sequence $x$)
  • if $x$ is empty, then $F(x)$ is the empty forest;
  • if $x[1\ldots m]$ is not empty, let $(r_{1},\ldots,r_{k})$ be the ordered sequence of all the $k$ positions of the smallest value of $x$ and let $r_{k+1}=m+1$, then $F(x)$ is a forest composed of $k$ trees whose roots are $r_{1},\ldots,r_{k}$, such that:
    • the left sub-forest of $r_{1}$ is the Cartesian Forest $F(x[1\ldots r_{1}-1])$, and
    • for $1\leq i\leq k$, the right sub-forest of $r_{i}$ is the Cartesian Forest $F(x[r_{i}+1\ldots r_{i+1}-1])$.
Figure 2: The Cartesian Forest associated to an ordered sequence $x$. Roots $r_{2}$ to $r_{4}$ are created since the sequence contains a value equal to the one at position 55. The values between two roots $r_{i}$ and $r_{i+1}$ belong to the tree enrooted in $r_{i}$. Therefore, except for the leftmost tree of a Cartesian Forest, a tree cannot have a left sub-forest. This idea is true at every level and can be seen in this example at positions $4$ and $12$.
4 Exact Cartesian Forest Matching
We adapt the linear representations of Cartesian Trees, such as the Parent-Distance [19] and the Skipped-Number representation [6] to Cartesian Forests. Basically, for a sequence $x$, when computing $C_{i}(x)$ from $C_{i-1}(x)$, the Parent-Distance of $i$ is equal to the distance between $i$ and its parent in $C_{i}(x)$ and its Skipped-Number is the number of nodes removed from the right branch of $C_{i-1}(x)$ comparing to the right branch of $C_{i}(x)$ (see [3]).
The main idea of the adaptation to Cartesian Forests is simple: both representations involve the comparisons of two values $x[j]$ and $x[k]$ in the original sequence $x$, for any $j<k$. Suppose we have either $x[j]>x[k]$ or $x[j]<x[k]$, then we are in the case of Cartesian Trees and the linear representation is the same. The idea for the Parent-Distance of the Cartesian Forest of a sequence $x$ is that its absolute value at a position $i$ gives the distance between a node $i$ and its parent or left sibling in the Cartesian Forest associated to the sequence $x[1\ldots i]$. Let $\operatorname{equal}_{x}(i)=\max(\{k\mid x[k]=x[i]\mbox{ with }1\leq k<i\}\cup\{0\})$.
Definition 6(Forest Parent-Distance representation $\operatorname{F-PD}_{x}$)
Given a sequence $x$ of length $m$, the Forest Parent-Distance representation of $x$ is an integer sequence $\operatorname{F-PD}_{x}[1\ldots m]$ defined as follows:
$\operatorname{F-PD}_{x}[i]=\begin{cases}i-\operatorname{small}_{x}(i)&\text{if }\operatorname{small}_{x}(i)>\operatorname{equal}_{x}(i)\cr-(i-\operatorname{equal}_{x}(i))&\text{if }\operatorname{small}_{x}(i)<\operatorname{equal}_{x}(i)\cr 0&\text{otherwise.}\end{cases}$
The referent table that we define next contains the position of the next smaller (or equal) value in the sequence.
Definition 7(The referent table $\operatorname{F-ref}_{x}$)
Given a sequence $x[1\ldots m]$, the referent table of $x$ is a sequence of sets $\operatorname{F-ref}_{x}[1\ldots m]$ such that
$\operatorname{F-ref}_{x}[i]=\begin{cases}\min_{i<j\leq m}\{j\mid x[j]\leq x[i]\}&\text{if such }j\text{ exists}\cr-1&\text{ otherwise.}\end{cases}$
Definition 8(Forest Skipped-Number representation $\operatorname{F-SN}_{x}$)
Given a sequence $x[1\ldots m]$, the Forest Skipped-Number representation of $x$ is an integer sequence $\operatorname{F-SN}_{x}[1\ldots m]$ such that
$\operatorname{F-SN}_{x}[i]=\begin{cases}|\{j<i\mid\operatorname{F-ref}_{x}[j]=i\}|&\mbox{if }\operatorname{small}_{x}(i)>\operatorname{equal}_{x}(i)\cr-|\{j<i\mid\operatorname{F-ref}_{x}[j]=i\}|&\mbox{if }\operatorname{small}_{x}(i)<\operatorname{equal}_{x}(i)\cr\end{cases}$
Figure 3 shows examples of Forest Parent-Distance representations, Forest Skipped-Number representations and referent tables of two Cartesian Forests.
Figure 3: Two sequences $x$ and $y$, their associated Cartesian Forests $F(x)$ and $F(y)$, and their corresponding Forest Parent-Distance representations, Forest Skipped-Number representations and referent tables. As one can see, $x$ is a prefix of $y$ and the forest $F(x)$ is transformed into a sequence of left subtrees in $F(y)$.
Definition 9(Cartesian Forest Matching (CFM))
Given two sequences $p[1\ldots m]$ and $t[1\ldots n]$, find every position $j$, with $1\leq j\leq n-m+1$, such that $t[j\ldots j+m-1]\approx_{CF}p[1\ldots m]$.
Example 1
Let $t=(5,7,3,6,3,7,2,8,2,4,3,3)$ and $p=(2,3,1,4,1,5)$ respectively be the text and the pattern. We have two occurrences of $p$ in $t$ as $t[1\ldots 6]=(5,7,3,6,3,7)\approx_{CF}p$ and $t[5\ldots 10]=(3,7,2,8,2,4)\approx_{CF}p.$